OpenStudy (anonymous):

Use the quadratic formula to solve for x: x+1/x = x/2 I'll give you a medal ;)

OpenStudy (anonymous):

\[- (1/2) x = 1/x\] \[-(1/2) x^{2} -1 = 0\] \[x = \frac{ -b +/- \sqrt{b^{2} - 4ac} }{ 2a }\] \[x= \frac{ 0 +/- \sqrt{0 - 2}}{ -1 }\]

OpenStudy (anonymous):

the answer in the book is \[ 1+/-\sqrt{3}\]

OpenStudy (anonymous):

but I don't know how they got that -.- and the answer you've got is totally different...

OpenStudy (anonymous):

is the equation \[x + \frac{ 1 }{ x } = \frac{ x }{ 2 }\]

OpenStudy (anonymous):

No it's \[\frac{ x+1 }{ x } = \frac{ x }{ 2 }\]

OpenStudy (anonymous):

\[x*x/(x+1) = 2\] \[x^{2}=2x+2\] \[x^{2}-2x-2 = 0\]

OpenStudy (phi):

to do this you first put the equation in standard form: \[ \frac{x+1}{x}= \frac{x}{2}\] you put it in standard form by getting rid of the fractions. multiply both sides by x \[ x\cdot \frac{x+1}{x}= x \cdot \frac{x}{2}\] which simplifies \[ \cancel{x}\cdot \frac{x+1}{\cancel{x}}=\frac{x^2}{2}\] now multiply both sides by 2 \[ 2(x+1)= 2 \cdot \frac{x^2}{2}\] cancel the 2's on the right side \[ 2(x+1)= \cancel{2} \cdot \frac{x^2}{\cancel{2}}\] on the left side, distribute the 2 by multipying by each term inside the parens \[ 2x+2 = x^2\] add -2x -2 to both sides \[ 2x+2 -2x-2= x^2-2x-2\] finally we get \[ x^2-2x-2 =0\] now apply the quadratic formula see http://www.khanacademy.org/math/algebra/quadratics/v/applying-the-quadratic-formula